# Sample Compression Unleashed: New Generalization Bounds for Real Valued Losses

### v2

@@ 1 Introduction
Sample compression theory, introduced by Littlestone and Warmuth, (1986), is based on the fundamental idea that “compressing implies learning” (David et al.,, 2016). If it is possible to provably show that a learned model can be completely defined by a subset of the training dataset, then sample compression theory gives us generalization guarantees. The most well-known learning algorithms that comply with the sample compression framework are the support vector machine (Boser et al.,, 1992) and the perceptron (Rosenblatt,, 1958; Moran et al.,, 2020); the relevant training subset being formed by the support vectors in the former case, and the points causing an update of the predictor in the latter case. More recently, Snyder and Vishwanath, (2020) and Paccagnan et al., (2024) have introduced the first sample compression results for neural networks.
@@
The sample compression theory is rich and multiple different approaches exist. For example, Attias et al., (2018); Ben-David et al., (2024); David et al., (2016); Floyd and Warmuth, (1995); Hanneke and Kontorovich, (2021); Hanneke et al., (2018, 2019, 2024); Moran and Yehudayoff, (2016); Rubinstein and Rubinstein, (2012) propose theoretical results relating the VC dimension (Vapnik and Chervonenkis,, 1971) and the compression analysis. By relating the probability of change of compression to the true risk, Campi and Garatti, (2023); Paccagnan et al., (2024) express very tight guarantees for the consistent case, i.e., when the error on the training set is zero. Finally, Laviolette et al., (2005); Marchand et al., (2003); Marchand and Shawe-Taylor, (2002); Marchand and Sokolova, (2005); Shah, (2007) give computable risk certificates valid even in the non-consistent case.
@@
In this paper, we build on the setting of Laviolette et al., (2005), based on the binomial test-set bound of Langford, (2005), which by definition is the tightest bound for the zero-one loss under the sole i.i.d. assumption. However, the use of the zero-one loss restricts its application to supervised classification problems. By leveraging proof techniques from the PAC-Bayesian literature, we extend the framework to real-valued losses and open the way to obtaining bounds directly for the cross-entropy loss (Pérez-Ortiz et al.,, 2021) and unbounded losses (Haddouche et al.,, 2021; Casado et al.,, 2024; Rodríguez-Gálvez et al.,, 2024), for example under the sub-Gaussian assumption. Finally, we train deep neural networks and random forests with the Pick-To-Learn meta-algorithm (Paccagnan et al.,, 2024), an algorithm that modifies the training loop of a model to yield a sample-compressed predictor, and assess the tightness of our bounds in different settings. In the consistent case, our bounds are arbitrarily tight upper bounds on previous results restricted to the zero-one loss.
@@
Of note, a major asset of our sample-compress bounds is that they do not depend on the number of learnable parameters. Two models of different sizes can achieve the same guarantees as long as they achieve the same empirical loss using the same amount of data. This lets us train large models such as DistilBERT and still achieve tight generalization bounds.
@@ 2 Background and Notation
We are interested in the supervised learning framework. Let $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$ be a dataset of $n$ datapoints, with each point sampled i.i.d. (independently and identically distributed) from an unknown distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathbb{R}}^{d}\times\operatorname{\mathcal{Y}}$. The targets are defined by the task at hand, with $\operatorname{\mathcal{Y}}\in\{-1,+1\}$ for binary classification tasks and $\operatorname{\mathcal{Y}}\subseteq\operatorname{\mathbb{R}}$ for regression tasks. For the rest of this section, we focus on binary classification problems, but in Section 3, we study both classification and regression settings.
Let $\operatorname{\mathcal{H}}$ be a family of predictors $h:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$. Let $A:\bigcup_{k=1}^{\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{k}\to\operatorname{\mathcal{H}}$ be an algorithm that takes a dataset $S$ and returns a predictor $A(S)$. We consider the zero-one loss function $\ell^{0\textrm{-}1}(h,\boldsymbol{x},y)=\operatorname{\mathbb{I}}[h(\boldsymbol{x})\neq y]$, with $\operatorname{\mathbb{I}}[a]=1$ if the predicate $a$ is true and 00 otherwise. Then, the true risk of the hypothesis $h$ is defined as
@@ 2.1 Sample compression theory
Let $h=A(S)$ be the output of algorithm $A$ applied to a dataset $S$. In order to obtain guarantees on the generalization performance of $h$ using the sample compression theory, we need to be able to uniquely define $h$ as a function (the reconstruction function) of a subset of $S$ (the compression set) and a complementary sequence of information (the message).
The compression set $S_{\operatorname{\mathbf{i}}}$ is defined using a vector of indices $\operatorname{\mathbf{i}}=\quantity(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})$, where the indices are ordered such that $1\leq i_{1}<i_{2}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n$. The vector $\operatorname{\mathbf{i}}$ belongs in the set of all possible vectors composed of the natural numbers 1 through $n$, denoted $\operatorname{\mathscr{P}}(n)$. Using this notation, $\operatorname{\mathbf{i}}$ indicates the datapoints of $S$ that are present in $S_{\operatorname{\mathbf{i}}}$, as such
@@
The message $\sigma$ is chosen in a set $M(\operatorname{\mathbf{i}})$, which contains all relevant messages associated to the compression set $\operatorname{\mathbf{i}}$. The message is a complementary source of information and is generally defined as a binary sequence.
@@
We define a distribution $P_{\overline{\operatorname{\mathcal{H}}}}$ over $\overline{\operatorname{\mathcal{H}}}$, such that $\sum_{h\in\overline{\operatorname{\mathcal{H}}}}P_{\overline{\operatorname{\mathcal{H}}}}(h)\leq 1$. As all sample-compressed predictors are uniquely defined using the indices vector and the message, we choose the distribution $P_{\overline{\operatorname{\mathcal{H}}}}$ to be a product of two distributions $P_{\overline{\operatorname{\mathcal{H}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{M(\operatorname{\mathbf{i}})}(\sigma)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a distribution on $\operatorname{\mathscr{P}}(n)$ and $P_{M(\operatorname{\mathbf{i}})}$ a distribution on $M(\operatorname{\mathbf{i}})$. Following previous works (e.g. Marchand and Sokolova,, 2005), we require the distribution $P_{\operatorname{\mathcal{H}}}$ to be data-independent, in order to avoid further assumptions. Without any information on the data, we generally set $P_{M(\operatorname{\mathbf{i}})}$ to a uniform distribution. As for the distribution $P_{\operatorname{\mathscr{P}}(n)}$, it is usually set to penalize larger compression sets (Laviolette et al.,, 2005; Marchand and Shawe-Taylor,, 2002; Marchand and Sokolova,, 2005). For any size of compression set $\operatorname{|\mathbf{i}|}$, there are $\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})$ different possible compression sets. We set the distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})$ to be $\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(m)=\tfrac{6}{\pi^{2}}(m+1)^{-2}$. This choice is discussed by Marchand and Sokolova, (2005).
@@
We now present the sample compression bound of Laviolette et al., (2005). This result is derived using the binomial test-set bound of Langford, (2005), which by definition is the tightest bound for the zero-one loss under the sole i.i.d. assumption.
@@
$R_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq\overline{\emph{Bin}}\quantity(\kappa,n,\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta)$
with $\kappa=n\widehat{R}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))$ and
@@
This theorem can be applied to any family of sample-compressed predictors, such as the support vector machine, the perceptron, and the set covering machine (Marchand and Shawe-Taylor,, 2002). To apply this theorem to neural networks, one must design a reconstruction function outputting neural networks. To this end, Snyder and Vishwanath, (2020) propose to reparameterize a 2-layer LeakyReLU network in order to obtain “support vectors”, which become the compression set of the reconstructed network. The following section presents a more general approach proposed by Paccagnan et al., (2024).
@@ 2.2 Pick-To-Learn
Initialize : $S_{\operatorname{\mathbf{i}}}=\emptyset$
Initialize : $h_{\operatorname{\mathbf{i}}}=h_{0}$
Initialize : $(\overline{\boldsymbol{x}},\overline{y})=\operatorname*{arg\,max}_{(\boldsymbol{x},y)\in S}\ell^{\textrm{x-e}}(h_{0},\boldsymbol{x},y)$
@@
with
$\Psi_{k,\delta}(\varepsilon)$ $=\frac{\delta}{2N}\sum_{m=k}^{n-1}\frac{\smallmatrixquantity(m\\ k)}{\smallmatrixquantity(n\\ k)}(1-\varepsilon)^{-(n-m)}$
$+\frac{\delta}{6N}\sum_{m=n+1}^{4N}\frac{\smallmatrixquantity(m\\ k)}{\smallmatrixquantity(n\\ k)}(1-\varepsilon)^{m-n}$
and where, for $k=0,1,\ldots,n-1$, $\overline{\varepsilon}(k,\delta)$ is the unique solution to the equation $\Psi_{k,\delta}(\varepsilon)=1$ in the interval $[\frac{k}{n},1]$, while $\overline{\varepsilon}(n,\delta)=1$.
Note that the value of previous bound is completely determined by the size of the compression set. The faster P2L obtains zero errors, the better the bound will be.
3 A General Sample-Compress Bound

### v3

@@ 1 INTRODUCTION
Sample compression theory, introduced by Littlestone and Warmuth, (1986), is based on the fundamental idea that “compressing implies learning” (David et al.,, 2016). If it is possible to provably show that a learned model can be completely defined by a subset of the training dataset, then sample compression theory gives us generalization guarantees. The most well-known learning algorithms that comply with the sample compression framework are the support vector machine (SVM) (Boser et al.,, 1992) and the perceptron (Rosenblatt,, 1958; Moran et al.,, 2020); the relevant training subset being formed by the support vectors in the former case, and the points causing an update of the predictor in the latter case. More recently, Snyder and Vishwanath, (2020) and Paccagnan et al., (2024) have introduced the first sample compression results for neural networks.
@@
The sample compression theory is rich and multiple different approaches exist. For example, Attias et al., (2018, 2023, 2024); Ben-David et al., (2024); David et al., (2016); Floyd and Warmuth, (1995); Hanneke and Kontorovich, (2021); Hanneke et al., (2018, 2019, 2024); Moran and Yehudayoff, (2016); Rubinstein and Rubinstein, (2012) propose theoretical results relating the VC dimension (Vapnik and Chervonenkis,, 1971) and the compression analysis. By relating the probability of change of compression to the true risk, Campi and Garatti, (2023); Paccagnan et al., (2024) express very tight guarantees for the consistent case, i.e., when the error on the training set is zero. Finally, Marchand and Shawe-Taylor, (2002); Marchand et al., (2003); Graepel et al., (2005); Laviolette et al., (2005, 2009); Marchand and Sokolova, (2005); Hussain et al., (2007); Shah, (2007) give computable risk certificates valid even in the non-consistent case.
@@
In this paper, we build on the setting of Laviolette et al., (2005). Their sample-compression bound is based on the binomial test-set bound of Langford, (2005), which by definition is the tightest test-set bound for the zero-one loss under the sole i.i.d. assumption. However, the use of the zero-one loss restricts its application to supervised classification problems. By leveraging proof techniques from the PAC-Bayesian literature, we extend the framework to real-valued losses and open the way to obtaining bounds directly for the cross-entropy loss (Pérez-Ortiz et al.,, 2021) and unbounded losses (Haddouche et al.,, 2021; Casado Telletxea et al.,, 2025; Rodríguez-Gálvez et al.,, 2024), for example under the sub-Gaussian assumption (Kahane,, 1960). Finally, we train deep neural networks and random forests with Pick-To-Learn (P2L) (Paccagnan et al.,, 2024), a meta-algorithm that modifies the training loop of a model to yield a sample-compressed predictor, and assess the tightness of our bounds in different settings.
@@
Of note, a major asset of our sample-compress bounds is that they do not depend on the number of learnable parameters. Two models of different sizes can achieve the same guarantees as long as they achieve the same empirical loss using the same amount of data. This lets us train large models such as DistilBERT (Sanh et al.,, 2019) and still achieve tight generalization bounds.
The paper is organized as follows. In Section 2, we present the sample compression theory and the meta-algorithm Pick-To-Learn (P2L) (Paccagnan et al.,, 2024). In Section 3, we first present Theorem 3, a new general sample-compression theorem that holds for any real-valued losses. Leveraging the comparator functions of PAC-Bayes theory (McAllester,, 1998), we present two new sample compression bounds for losses in the interval unit, Corollary 4 and Corollary 6, which respectively yield the tightest bound in theory and in practice. Then, we present Corollary 7, which holds for any unbounded losses, under the assumption that the moment-generating function is bounded. We finish this section by proving the tightness of our results over the previous state-of-the-art sample compression bound. Finally, in Section 4, we empirically show the tightness of our bounds by training deep neural networks on image and text classification problems with P2L. We adapt P2L to regression, train regression trees and forests with this modified algorithm and provide the first sample compression generalization bounds for tree-based regression predictors.
@@ 2 BACKGROUND AND NOTATION
We are interested in the supervised learning framework. Let $(\boldsymbol{x}_{1},y_{1}),\ldots,(\boldsymbol{x}_{n},y_{n})$ be a sequence of $n$ datapoints sampled i.i.d. (independently and identically distributed) from an unknown distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathbb{R}}^{d}\times\operatorname{\mathcal{Y}}$. The dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$ is generated with the sequence of datapoints. [^1] In this paper, we do not consider repeated datapoints. However, all definitions and results could be easily adapted to use multisets to account for repetitions, similarly to the work of Campi and Garatti, (2023). The targets are defined by the task at hand, with $\operatorname{\mathcal{Y}}\in\{-1,+1\}$ for binary classification tasks and $\operatorname{\mathcal{Y}}\subseteq\operatorname{\mathbb{R}}$ for regression tasks. In this section, we focus on binary classification problems, but in Section 3, we study both classification and regression settings.
Let $\operatorname{\mathcal{H}}$ be a family of predictors $h:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$. Let $A:\bigcup_{k=1}^{\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{k}\to\operatorname{\mathcal{H}}$ be a learning algorithm that takes a dataset $S$ and returns a predictor $A(S)$. We consider the zero-one loss function $\ell^{0\textrm{-}1}(h,\boldsymbol{x},y)=\operatorname{\mathbb{I}}[h(\boldsymbol{x})\neq y]$, with $\operatorname{\mathbb{I}}[a]=1$ if the predicate $a$ is true and 00 otherwise. Then, the true risk of the hypothesis $h$ is defined as
@@ 2.1 Sample compression theory
Let the predictor $h=A(S)$ be the output of a learning algorithm $A$ applied to a dataset $S$. In order to obtain guarantees on the generalization performance of $h$ using the sample compression theory, we need to be able to uniquely define $h$ as a function (the reconstruction function) of a subset of $S$ (the compression set) and a complementary sequence of information (the message).
The compression set $S_{\operatorname{\mathbf{i}}}$ is defined using a vector of indices $\operatorname{\mathbf{i}}=\quantity(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})$, where the indices are ordered such that $1\leq i_{1}<i_{2}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n$. The vector $\operatorname{\mathbf{i}}$ belongs in the set of all possible vectors composed of the natural numbers 1 through $n$, denoted
$\operatorname{\mathscr{P}}(n)=\bigg{\{}\emptyset,\{1\},\{2\},\ldots,\{n\},\{1,2\},\ldots,\{1,n\},\ldots,\{1,2,\ldots,n\}\bigg{\}}.$
Using this notation, $\operatorname{\mathbf{i}}$ indicates the datapoints of $S$ that are present in $S_{\operatorname{\mathbf{i}}}$ :
@@
The message $\sigma$ is chosen in a set $M(\operatorname{\mathbf{i}})$, which contains all relevant messages associated to the compression set $\operatorname{\mathbf{i}}$. The message is a complementary source of information needed to reconstruct the predictor.
@@
In this paper, we distinguish two categories of reconstruction functions: inherent and dedicated. An inherent reconstruction function is used when a learning algorithm $A$ is its own reconstruction function. An algorithm $A$ is an inherent reconstruction function when, given a dataset $S$ and its compression set $S_{\operatorname{\mathbf{i}}}$, the following equality holds $A(S)=A(S_{\operatorname{\mathbf{i}}})$. The most well-known example of an inherent reconstruction function is the SVM (Boser et al.,, 1992). Other examples of inherent reconstruction functions are the perceptron (Rosenblatt,, 1958) and Pick-To-Learn (Paccagnan et al.,, 2024). On the other hand, dedicated reconstruction functions are used when an algorithm cannot be used to reconstruct the learned predictor from a compression set. The dedicated reconstruction function is different from the learning algorithm $A$ and is generally hand-crafted to suit $A$. The reconstruction function of the SCM (Marchand and Shawe-Taylor,, 2002) and all its iterations (e.g., Marchand et al.,, 2003; Marchand and Sokolova,, 2005; Laviolette et al.,, 2005; Kestler et al.,, 2006; Hussain et al.,, 2007; Drouin et al.,, 2019) are examples of dedicated reconstruction functions.
We provide an example of a dedicated reconstruction function for a very simple predictor, the decision stump.
Example (Shah et al., (2011)).
Given a datapoint $\boldsymbol{x}^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{d})$, a direction {$,+1\}$ 1 } and an index $1\leq k\leq d$, the stump is defined $f_{(\boldsymbol{x}^{\prime},d,k)}(\boldsymbol{x})=\operatorname{\mathbb{I}}\quantity[\diamond\cdot(x_{k}-x^{\prime}_{k})>0]$. To learn a decision stump over a dataset $S$, each combination of $\boldsymbol{x}^{\prime}\in S$, {$,+1\}$ 1 } and $0\leq k\leq d$ is tested. Once the decision stump is learned, it is completely defined by the datapoint $\boldsymbol{x}^{\prime}$, the direction $\diamond$ and the index $k$. Our compression set is $S_{\operatorname{\mathbf{i}}}=\{\boldsymbol{x}^{\prime}\}$ and the message is $\sigma=\{\diamond,k\}$. With the compression set and the message, we can fully reconstruct the stump with $\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\{\diamond,k\})=f_{(S_{\operatorname{\mathbf{i}}},\diamond,k)}$.
@@
Let $P_{\overline{\operatorname{\mathcal{H}}}}$ be a distribution over $\overline{\operatorname{\mathcal{H}}}$, such that $\sum_{h\in\overline{\operatorname{\mathcal{H}}}}P_{\overline{\operatorname{\mathcal{H}}}}(h)\leq 1$. As all sample-compressed predictors are uniquely defined using the index vector and the message, we choose the distribution $P_{\overline{\operatorname{\mathcal{H}}}}$ to be a product of two distributions $P_{\overline{\operatorname{\mathcal{H}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{M(\operatorname{\mathbf{i}})}(\sigma)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a distribution on $\operatorname{\mathscr{P}}(n)$ and $P_{M(\operatorname{\mathbf{i}})}$ a distribution on $M(\operatorname{\mathbf{i}})$. Following previous works (e.g. Marchand and Sokolova,, 2005), we require the distribution $P_{\operatorname{\mathcal{H}}}$ to be data-independent, in order to avoid further assumptions. Without any information on the data, we generally set $P_{M(\operatorname{\mathbf{i}})}$ to a uniform distribution. As for the distribution $P_{\operatorname{\mathscr{P}}(n)}$, it is usually set to penalize larger compression sets (Laviolette et al.,, 2005; Marchand and Shawe-Taylor,, 2002; Marchand and Sokolova,, 2005). For any size of compression set $\operatorname{|\mathbf{i}|}$, there are $\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})$ different possible compression sets. We set the distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})$ to be $\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(m)=\tfrac{6}{\pi^{2}}(m+1)^{-2}$. This choice is discussed by Marchand and Sokolova, (2005).
@@
We now present the sample compression bound of Laviolette et al., (2005). This result is derived using the binomial test-set bound of Langford, (2005), which by definition is the tightest test-set bound for the zero-one loss under the sole i.i.d. assumption.
@@
$R_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq\overline{\emph{Bin}}\quantity(\kappa,|\operatorname{\mathbf{i}}^{c}|,\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta),$
with $\kappa=|\operatorname{\mathbf{i}}^{c}|\widehat{R}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))$ and
@@
This theorem can be applied to any family of sample-compressed predictors, such as the support vector machine (Boser et al.,, 1992), the perceptron (Rosenblatt,, 1958) and the set covering machine (Marchand and Shawe-Taylor,, 2002). To apply this theorem to neural networks, one must design a reconstruction function outputting neural networks. To this end, Snyder and Vishwanath, (2020) propose to reparameterize a 2-layer LeakyReLU network in order to obtain “support vectors”, which become the compression set of the reconstructed network. The following section presents a more general approach proposed by Paccagnan et al., (2024).
@@ 2.2 Pick-To-Learn
Initialize : $S_{\operatorname{\mathbf{i}}}\leftarrow\emptyset$
Initialize : $h_{\operatorname{\mathbf{i}}}\leftarrow h_{0}$
Initialize : $(\overline{\boldsymbol{x}},\overline{y})\leftarrow\operatorname*{arg\,max}_{(\boldsymbol{x},y)\in S}\ell^{\textrm{x-e}}(h_{0},\boldsymbol{x},y)$
@@
where, for $k=0,1,\ldots,n-1$, $\overline{\varepsilon}(k,\delta)$ is the unique solution to the equation $\Psi_{k,\delta}(\varepsilon)=1$ in the interval $[\frac{k}{n},1]$, with
$\Psi_{k,\delta}(\varepsilon)$ $=\frac{\delta}{2n}\hskip 4.2679pt\sum_{m=k}^{n-1}\ \ \frac{\smallmatrixquantity(m\\ k)}{\smallmatrixquantity(n\\ k)}(1-\varepsilon)^{-(n-m)}$
$+\frac{\delta}{6n}\sum_{m=n+1}^{4n}\frac{\smallmatrixquantity(m\\ k)}{\smallmatrixquantity(n\\ k)}(1-\varepsilon)^{-(n-m)}\,,$
and $\overline{\varepsilon}(n,\delta)=1$.
Note that the value of the previous bound is completely determined by $|\operatorname{\mathbf{i}}|$, the size of the compression set. The faster P2L obtains zero errors (in terms of the number of iterations performed by Algorithm 1), the better the bound.
3 A GENERAL SAMPLE-COMPRESSION BOUND